# Key Words.: Stars: evolution – Stars: general – Stars: massive – Stars: abundances – Stars: black holes – Methods: numerical1 Introduction

### v1

@@ Appendix F Extra tables
14 | 11.93 | 14 | CCSN | 1.49 | 3.84 | 2.79 | 0.01 | 9.35 | 1.44 |
20 | 8.08 | 20 | CCSN | 1.57 | 6.73 | 4.97 | 0.00 | 9.40 | 1.42 |
30 | 5.75 | 30 | FSN | 29.50 | 12.08 | 9.79 | 0.05 | 9.39 | 1.41 |
40 | 4.64 | 39.99 | FSN | 39.50 | 16.91 | 15.08 | 0.02 | 9.45 | 1.38 |
50 | 4.07 | 49.99 | FSN | 49.50 | 22.43 | 19.65 | 0.02 | 9.45 | 1.36 |
80 | 3.22 | 79.99 | FSN | 79.50 | 33.08 | 30.21 | 0.23 | 9.42 | 1.34 |
100 | 2.95 | 99.99 | FSN | 99.50 | 32.96 | 31.64 | 0.16 | 9.43 | 1.34 |
200 | 2.37 | 199.99 | PISN | - | 110.08 | 106.15 | 0.80 | 9.19 | 1.33 |
300 | 2.14 | 299.99 | DBH | 299.50 | 169.26 | 164.48 | 0.79 | 9.14 | 1.33 |
400 | 2.04 | 399.99 | DBH | 399.49 | 227.87 | 221.41 | 0.77 | 9.11 | 1.33 |
500 | 1.94 | 499.98 | DBH | 499.48 | 286.85 | 277.39 | 0.75 | 9.09 | 1.33 |
600 | 1.90 | 599.97 | DBH | 599.47 | 345.96 | 333.83 | 0.74 | 9.08 | 1.33 |
1000 | 1.80 | 999.81 | DBH | 9993.1 | 576.94 | 538.72 | 0.70 | 9.02 | 1.33 |
2000 | 1.65 | 1997.26 | DBH | 1996.76 | 1174.06 | 1103.56 | 0.61 | 8.98 | 1.33 |
14 | 14.09 | 13.99 | CCSN | 1.49 | 4.54 | 2.79 | 0.00 | 9.35 | 1.41 |
20 | 9.08 | 19.99 | CCSN | 1.59 | 7.32 | 5.15 | 0.06 | 9.38 | 1.42 |
30 | 6.16 | 29.99 | FSN | 29.49 | 12.75 | 10.09 | 0.00 | 9.41 | 1.41 |
40 | 4.91 | 39.99 | FSN | 39.49 | 18.15 | 15.25 | 0.00 | 9.44 | 1.38 |
50 | 4.28 | 49.99 | FSN | 49.49 | 22.96 | 21.04 | 0.08 | 9.43 | 1.37 |
80 | 3.37 | 79.98 | FSN | 79.48 | 28.84 | 26.32 | 0.04 | 9.40 | 1.35 |
100 | 3.06 | 99.96 | PPISN | 43.10 | 49.92 | 44.12 | 0.66 | 9.36 | 1.33 |
200 | 2.46 | 199.93 | PISN | - | 111.07 | 103.03 | 0.80 | 9.18 | 1.33 |
300 | 2.24 | 299.73 | DBH | 299.23 | 173.02 | 164.90 | 0.79 | 9.13 | 1.33 |
400 | 2.12 | 399.24 | DBH | 398.74 | 234.43 | 223.38 | 0.78 | 9.10 | 1.33 |
500 | 2.05 | 498.48 | DBH | 497.98 | 294.70 | 283.18 | 0.76 | 9.08 | 1.33 |
600 | 1.99 | 597.55 | DBH | 597.05 | 352.42 | 338.47 | 0.74 | 9.08 | 1.33 |
1000 | 1.88 | 989.94 | DBH | 989.45 | 593.16 | 567.88 | 0.69 | 9.04 | 1.33 |
2000 | 1.76 | 1950.40 | DBH | 1981.27 | 1169.90 | 1106.60 | 0.61 | 8.78 | 1.34 |
14 | 14.47 | 13.99 | CCSN | 1.49 | 4.91 | 2.84 | 0.00 | 9.35 | 1.43 |
20 | 9.29 | 19.97 | CCSN | 1.57 | 7.56 | 4.98 | 0.02 | 9.40 | 1.42 |
30 | 6.29 | 29.95 | FSN | 29.45 | 12.65 | 9.58 | 0.04 | 9.38 | 1.41 |
40 | 5.04 | 39.90 | FSN | 39.40 | 18.44 | 14.88 | 0.00 | 9.43 | 1.39 |
50 | 4.37 | 49.84 | FSN | 49.33 | 24.48 | 20.44 | 0.02 | 9.44 | 1.37 |
80 | 3.42 | 78.88 | PPISN | 38.69 | 42.83 | 37.68 | 0.53 | 9.39 | 1.34 |
100 | 3.15 | 97.18 | FSN | 96.68 | 32.78 | 32.45 | 0.06 | 9.39 | 1.35 |
200 | 2.51 | 196.45 | PISN | - | 113.17 | 104.09 | 0.80 | 9.18 | 1.33 |
300 | 2.29 | 291.76 | DBH | 291.26 | 174.97 | 160.07 | 0.79 | 9.12 | 1.33 |
400 | 2.17 | 384.07 | DBH | 383.57 | 231.80 | 212.58 | 0.78 | 9.10 | 1.33 |
500 | 2.09 | 473.87 | DBH | 473.38 | 296.34 | 273.56 | 0.76 | 9.08 | 1.33 |
600 | 2.01 | 560.44 | DBH | 559.94 | 359.77 | 336.05 | 0.74 | 9.06 | 1.33 |
1000 | 1.92 | 912.27 | DBH | 911.78 | 820.52 | 396.31 | 0.70 | 9.03 | 1.34 |
2000 | 1.75 | 1777.3 | DBH | 1776.89 | 553.51 | 553.18 | 0.55 | 9.11 | 1.33 |
@@
14 | 14.65 | 13.94 | CCSN | 1.42 | 4.94 | 2.83 | 0.01 | 9.33 | 1.43 |
20 | 9.40 | 19.82 | CCSN | 1.64 | 7.62 | 4.93 | 0.03 | 9.39 | 1.42 |
30 | 6.37 | 29.47 | FSN | 28.96 | 12.61 | 9.43 | 0.03 | 9.38 | 1.41 |
40 | 5.14 | 38.72 | FSN | 38.22 | 18.16 | 14.54 | 0.00 | 9.43 | 1.39 |
50 | 4.39 | 46.71 | FSN | 46.21 | 24.09 | 19.98 | 0.14 | 9.41 | 1.37 |
80 | 3.51 | 69.36 | PPISN | 38.78 | 30.59 | 27.65 | 0.04 | 9.39 | 1.35 |
100 | 3.19 | 75.08 | PPISN | 33.92 | 34.57 | 34.56 | 0.18 | 9.41 | 1.35 |
200 | 2.55 | 157.60 | PISN | - | 89.75 | 85.92 | 0.82 | 9.21 | 1.33 |
300 | 2.31 | 213.81 | DBH | 213.31 | 174.64 | 157.77 | 0.80 | 9.12 | 1.33 |
400 | 2.20 | 289.49 | DBH | 288.99 | 222.66 | 206.17 | 0.78 | 9.11 | 1.33 |
500 | 2.12 | 363.47 | DBH | 362.97 | 274.31 | 255.43 | 0.77 | 9.09 | 1.33 |
600 | 2.06 | 439.78 | DBH | 439.28 | 323.54 | 303.96 | 0.75 | 9.07 | 1.33 |
14 | 14.57 | 13.74 | CCSN | 1.37 | 4.78 | 2.69 | 0.01 | 9.36 | 1.43 |
20 | 9.31 | 18.08 | CCSN | 1.54 | 7.49 | 4.88 | 0.02 | 9.39 | 1.42 |
30 | 6.27 | 24.06 | FSN | 23.55 | 12.46 | 9.29 | 0.03 | 9.38 | 1.41 |
40 | 5.04 | 28.23 | FSN | 27.72 | 17.73 | 14.12 | 0.68 | 9.11 | 1.40 |
50 | 4.33 | 27.99 | FSN | 27.49 | 23.33 | 19.38 | 0.44 | 9.37 | 1.37 |
80 | 3.43 | 41.11 | PPISN | 38.70 | 40.28 | 35.82 | 0.85 | 9.26 | 1.35 |
100 | 3.17 | 52.08 | PPISN | 43.84 | 51.56 | 45.51 | 0.86 | 9.27 | 1.34 |
200 | 2.54 | 104.99 | PISN | - | 103.94 | 93.45 | 0.79 | 9.19 | 1.33 |
300 | 2.32 | 158.24 | DBH | 157.74 | 156.66 | 142.25 | 0.79 | 9.14 | 1.33 |
400 | 2.20 | 210.69 | DBH | 210.19 | 208.59 | 190.50 | 0.78 | 9.10 | 1.33 |
500 | 2.12 | 263.66 | DBH | 263.16 | 261.02 | 238.98 | 0.77 | 9.09 | 1.33 |
600 | 2.07 | 314.80 | DBH | 314.30 | 311.65 | 285.89 | 0.75 | 9.07 | 1.33 |
14 | 14.30 | 13.53 | CCSN | 1.42 | 4.76 | 2.71 | 0.01 | 9.36 | 1.43 |
20 | 9.15 | 18.52 | CCSN | 1.54 | 7.25 | 4.83 | 0.04 | 9.39 | 1.42 |
30 | 6.14 | 23.13 | FSN | 22.63 | 12.42 | 9.27 | 0.03 | 9.38 | 1.41 |
40 | 5.04 | 20.98 | FSN | 20.77 | 17.62 | 14.14 | 0.00 | 9.43 | 1.39 |
50 | 4.30 | 19.96 | FSN | 19.46 | 19.76 | 17.08 | 0.27 | 9.39 | 1.39 |
80 | 3.40 | 33.29 | FSN | 32.79 | 32.62 | 28.73 | 0.71 | 9.37 | 1.35 |
100 | 3.15 | 38.54 | PPISN | 37.53 | 38.15 | 33.69 | 0.61 | 9.38 | 1.34 |
200 | 2.55 | 73.60 | PISN | - | 72.87 | 64.99 | 0.85 | 9.24 | 1.33 |
300 | 2.33 | 108.34 | PISN | - | 107.26 | 96.79 | 0.79 | 9.18 | 1.33 |
400 | 2.22 | 141.49 | DBH | 140.99 | 140.07 | 127.20 | 0.79 | 9.15 | 1.33 |
500 | 2.14 | 172.54 | DBH | 172.05 | 170.82 | 155.66 | 0.78 | 9.12 | 1.33 |
600 | 2.10 | 186.83 | DBH | 186.33 | 184.96 | 168.80 | 0.78 | 9.12 | 1.33 |
14 | 13.86 | 13.10 | CCSN | 1.49 | 4.60 | 2.69 | 0.01 | 9.36 | 1.43 |
20 | 8.75 | 17.74 | CCSN | 1.55 | 7.10 | 4.85 | 0.68 | 9.20 | 1.43 |
30 | 5.91 | 13.21 | FSN | 12.70 | 12.81 | 9.32 | 0.05 | 9.38 | 1.41 |
40 | 4.90 | 8.93 | FSN | 8.43 | 8.93 | 7.23 | 0.00 | 9.42 | 1.40 |
50 | 4.21 | 11.84 | FSN | 11.33 | 11.59 | 9.64 | 0.17 | 9.37 | 1.39 |
80 | 3.39 | 15.86 | FSN | 15.36 | 15.61 | 13.09 | 0.11 | 9.40 | 1.39 |
100 | 3.18 | 17.38 | FSN | 16.88 | 17.21 | 14.74 | 0.04 | 9.41 | 1.38 |
200 | 2.65 | 25.05 | FSN | 24.55 | 24.80 | 21.02 | 0.50 | 9.38 | 1.36 |
300 | 2.44 | 33.19 | FSN | 32.69 | 32.86 | 28.63 | 0.61 | 9.38 | 1.35 |
400 | 2.35 | 34.97 | PPISN | 34.32 | 34.62 | 30.35 | 0.67 | 9.38 | 1.34 |
500 | 2.33 | 34.79 | PPISN | 34.14 | 34.45 | 30.14 | 0.68 | 9.37 | 1.34 |
600 | 2.32 | 34.77 | PPISN | 34.12 | 34.42 | 30.15 | 0.63 | 9.38 | 1.34 |

### v2

@@ Appendix F Extra tables
| 14 | 11.93 | 14 | CCSN | 1.49 | 3.84 | 2.79 | 0.01 | 9.35 | 1.44 |
| 20 | 8.08 | 20 | CCSN | 1.57 | 6.73 | 4.97 | 0.00 | 9.40 | 1.42 |
| 30 | 5.75 | 30 | FSN | 29.50 | 12.08 | 9.79 | 0.05 | 9.39 | 1.41 |
| 40 | 4.64 | 39.99 | FSN | 39.50 | 16.91 | 15.08 | 0.02 | 9.45 | 1.38 |
| 50 | 4.07 | 49.99 | FSN | 49.50 | 22.43 | 19.65 | 0.02 | 9.45 | 1.36 |
| 80 | 3.22 | 79.99 | FSN | 79.50 | 33.08 | 30.21 | 0.23 | 9.42 | 1.34 |
| 100 | 2.95 | 99.99 | FSN | 99.50 | 32.96 | 31.64 | 0.16 | 9.43 | 1.34 |
| 200 | 2.37 | 199.99 | PISN | - | 110.08 | 106.15 | 0.80 | 9.19 | 1.33 |
| 300 | 2.14 | 299.99 | DBH | 299.50 | 169.26 | 164.48 | 0.79 | 9.14 | 1.33 |
| 400 | 2.04 | 399.99 | DBH | 399.49 | 227.87 | 221.41 | 0.77 | 9.11 | 1.33 |
| 500 | 1.94 | 499.98 | DBH | 499.48 | 286.85 | 277.39 | 0.75 | 9.09 | 1.33 |
| 600 | 1.90 | 599.97 | DBH | 599.47 | 345.96 | 333.83 | 0.74 | 9.08 | 1.33 |
| 1000 | 1.80 | 999.81 | DBH | 9993.1 | 576.94 | 538.72 | 0.70 | 9.02 | 1.33 |
| 2000 | 1.65 | 1997.26 | DBH | 1996.76 | 1174.06 | 1103.56 | 0.61 | 8.98 | 1.33 |
| 14 | 14.09 | 13.99 | CCSN | 1.49 | 4.54 | 2.79 | 0.00 | 9.35 | 1.41 |
| 20 | 9.08 | 19.99 | CCSN | 1.59 | 7.32 | 5.15 | 0.06 | 9.38 | 1.42 |
| 30 | 6.16 | 29.99 | FSN | 29.49 | 12.75 | 10.09 | 0.00 | 9.41 | 1.41 |
| 40 | 4.91 | 39.99 | FSN | 39.49 | 18.15 | 15.25 | 0.00 | 9.44 | 1.38 |
| 50 | 4.28 | 49.99 | FSN | 49.49 | 22.96 | 21.04 | 0.08 | 9.43 | 1.37 |
| 80 | 3.37 | 79.98 | FSN | 79.48 | 28.84 | 26.32 | 0.04 | 9.40 | 1.35 |
| 100 | 3.06 | 99.96 | PPISN | 43.10 | 49.92 | 44.12 | 0.66 | 9.36 | 1.33 |
| 200 | 2.46 | 199.93 | PISN | - | 111.07 | 103.03 | 0.80 | 9.18 | 1.33 |
| 300 | 2.24 | 299.73 | DBH | 299.23 | 173.02 | 164.90 | 0.79 | 9.13 | 1.33 |
| 400 | 2.12 | 399.24 | DBH | 398.74 | 234.43 | 223.38 | 0.78 | 9.10 | 1.33 |
| 500 | 2.05 | 498.48 | DBH | 497.98 | 294.70 | 283.18 | 0.76 | 9.08 | 1.33 |
| 600 | 1.99 | 597.55 | DBH | 597.05 | 352.42 | 338.47 | 0.74 | 9.08 | 1.33 |
| 1000 | 1.88 | 989.94 | DBH | 989.45 | 593.16 | 567.88 | 0.69 | 9.04 | 1.33 |
| 2000 | 1.76 | 1950.40 | DBH | 1981.27 | 1169.90 | 1106.60 | 0.61 | 8.78 | 1.34 |
| 14 | 14.47 | 13.99 | CCSN | 1.49 | 4.91 | 2.84 | 0.00 | 9.35 | 1.43 |
| 20 | 9.29 | 19.97 | CCSN | 1.57 | 7.56 | 4.98 | 0.02 | 9.40 | 1.42 |
| 30 | 6.29 | 29.95 | FSN | 29.45 | 12.65 | 9.58 | 0.04 | 9.38 | 1.41 |
| 40 | 5.04 | 39.90 | FSN | 39.40 | 18.44 | 14.88 | 0.00 | 9.43 | 1.39 |
| 50 | 4.37 | 49.84 | FSN | 49.33 | 24.48 | 20.44 | 0.02 | 9.44 | 1.37 |
| 80 | 3.42 | 78.88 | PPISN | 38.69 | 42.83 | 37.68 | 0.53 | 9.39 | 1.34 |
| 100 | 3.15 | 97.18 | FSN | 96.68 | 32.78 | 32.45 | 0.06 | 9.39 | 1.35 |
| 200 | 2.51 | 196.45 | PISN | - | 113.17 | 104.09 | 0.80 | 9.18 | 1.33 |
| 300 | 2.29 | 291.76 | DBH | 291.26 | 174.97 | 160.07 | 0.79 | 9.12 | 1.33 |
| 400 | 2.17 | 384.07 | DBH | 383.57 | 231.80 | 212.58 | 0.78 | 9.10 | 1.33 |
| 500 | 2.09 | 473.87 | DBH | 473.38 | 296.34 | 273.56 | 0.76 | 9.08 | 1.33 |
| 600 | 2.01 | 560.44 | DBH | 559.94 | 359.77 | 336.05 | 0.74 | 9.06 | 1.33 |
| 1000 | 1.92 | 912.27 | DBH | 911.78 | 820.52 | 396.31 | 0.70 | 9.03 | 1.34 |
| 2000 | 1.75 | 1777.3 | DBH | 1776.89 | 553.51 | 553.18 | 0.55 | 9.11 | 1.33 |
@@
| 14 | 14.65 | 13.94 | CCSN | 1.42 | 4.94 | 2.83 | 0.01 | 9.33 | 1.43 |
| 20 | 9.40 | 19.82 | CCSN | 1.64 | 7.62 | 4.93 | 0.03 | 9.39 | 1.42 |
| 30 | 6.37 | 29.47 | FSN | 28.96 | 12.61 | 9.43 | 0.03 | 9.38 | 1.41 |
| 40 | 5.14 | 38.72 | FSN | 38.22 | 18.16 | 14.54 | 0.00 | 9.43 | 1.39 |
| 50 | 4.39 | 46.71 | FSN | 46.21 | 24.09 | 19.98 | 0.14 | 9.41 | 1.37 |
| 80 | 3.51 | 69.36 | PPISN | 38.78 | 30.59 | 27.65 | 0.04 | 9.39 | 1.35 |
| 100 | 3.19 | 75.08 | PPISN | 33.92 | 34.57 | 34.56 | 0.18 | 9.41 | 1.35 |
| 200 | 2.55 | 157.60 | PISN | - | 89.75 | 85.92 | 0.82 | 9.21 | 1.33 |
| 300 | 2.31 | 213.81 | DBH | 213.31 | 174.64 | 157.77 | 0.80 | 9.12 | 1.33 |
| 400 | 2.20 | 289.49 | DBH | 288.99 | 222.66 | 206.17 | 0.78 | 9.11 | 1.33 |
| 500 | 2.12 | 363.47 | DBH | 362.97 | 274.31 | 255.43 | 0.77 | 9.09 | 1.33 |
| 600 | 2.06 | 439.78 | DBH | 439.28 | 323.54 | 303.96 | 0.75 | 9.07 | 1.33 |
| 14 | 14.57 | 13.74 | CCSN | 1.37 | 4.78 | 2.69 | 0.01 | 9.36 | 1.43 |
| 20 | 9.31 | 18.08 | CCSN | 1.54 | 7.49 | 4.88 | 0.02 | 9.39 | 1.42 |
| 30 | 6.27 | 24.06 | FSN | 23.55 | 12.46 | 9.29 | 0.03 | 9.38 | 1.41 |
| 40 | 5.04 | 28.23 | FSN | 27.72 | 17.73 | 14.12 | 0.68 | 9.11 | 1.40 |
| 50 | 4.33 | 27.99 | FSN | 27.49 | 23.33 | 19.38 | 0.44 | 9.37 | 1.37 |
| 80 | 3.43 | 41.11 | PPISN | 38.70 | 40.28 | 35.82 | 0.85 | 9.26 | 1.35 |
| 100 | 3.17 | 52.08 | PPISN | 43.84 | 51.56 | 45.51 | 0.86 | 9.27 | 1.34 |
| 200 | 2.54 | 104.99 | PISN | - | 103.94 | 93.45 | 0.79 | 9.19 | 1.33 |
| 300 | 2.32 | 158.24 | DBH | 157.74 | 156.66 | 142.25 | 0.79 | 9.14 | 1.33 |
| 400 | 2.20 | 210.69 | DBH | 210.19 | 208.59 | 190.50 | 0.78 | 9.10 | 1.33 |
| 500 | 2.12 | 263.66 | DBH | 263.16 | 261.02 | 238.98 | 0.77 | 9.09 | 1.33 |
| 600 | 2.07 | 314.80 | DBH | 314.30 | 311.65 | 285.89 | 0.75 | 9.07 | 1.33 |
| 14 | 14.30 | 13.53 | CCSN | 1.42 | 4.76 | 2.71 | 0.01 | 9.36 | 1.43 |
| 20 | 9.15 | 18.52 | CCSN | 1.54 | 7.25 | 4.83 | 0.04 | 9.39 | 1.42 |
| 30 | 6.14 | 23.13 | FSN | 22.63 | 12.42 | 9.27 | 0.03 | 9.38 | 1.41 |
| 40 | 5.04 | 20.98 | FSN | 20.77 | 17.62 | 14.14 | 0.00 | 9.43 | 1.39 |
| 50 | 4.30 | 19.96 | FSN | 19.46 | 19.76 | 17.08 | 0.27 | 9.39 | 1.39 |
| 80 | 3.40 | 33.29 | FSN | 32.79 | 32.62 | 28.73 | 0.71 | 9.37 | 1.35 |
| 100 | 3.15 | 38.54 | PPISN | 37.53 | 38.15 | 33.69 | 0.61 | 9.38 | 1.34 |
| 200 | 2.55 | 73.60 | PISN | - | 72.87 | 64.99 | 0.85 | 9.24 | 1.33 |
| 300 | 2.33 | 108.34 | PISN | - | 107.26 | 96.79 | 0.79 | 9.18 | 1.33 |
| 400 | 2.22 | 141.49 | DBH | 140.99 | 140.07 | 127.20 | 0.79 | 9.15 | 1.33 |
| 500 | 2.14 | 172.54 | DBH | 172.05 | 170.82 | 155.66 | 0.78 | 9.12 | 1.33 |
| 600 | 2.10 | 186.83 | DBH | 186.33 | 184.96 | 168.80 | 0.78 | 9.12 | 1.33 |
| 14 | 13.86 | 13.10 | CCSN | 1.49 | 4.60 | 2.69 | 0.01 | 9.36 | 1.43 |
| 20 | 8.75 | 17.74 | CCSN | 1.55 | 7.10 | 4.85 | 0.68 | 9.20 | 1.43 |
| 30 | 5.91 | 13.21 | FSN | 12.70 | 12.81 | 9.32 | 0.05 | 9.38 | 1.41 |
| 40 | 4.90 | 8.93 | FSN | 8.43 | 8.93 | 7.23 | 0.00 | 9.42 | 1.40 |
| 50 | 4.21 | 11.84 | FSN | 11.33 | 11.59 | 9.64 | 0.17 | 9.37 | 1.39 |
| 80 | 3.39 | 15.86 | FSN | 15.36 | 15.61 | 13.09 | 0.11 | 9.40 | 1.39 |
| 100 | 3.18 | 17.38 | FSN | 16.88 | 17.21 | 14.74 | 0.04 | 9.41 | 1.38 |
| 200 | 2.65 | 25.05 | FSN | 24.55 | 24.80 | 21.02 | 0.50 | 9.38 | 1.36 |
| 300 | 2.44 | 33.19 | FSN | 32.69 | 32.86 | 28.63 | 0.61 | 9.38 | 1.35 |
| 400 | 2.35 | 34.97 | PPISN | 34.32 | 34.62 | 30.35 | 0.67 | 9.38 | 1.34 |
| 500 | 2.33 | 34.79 | PPISN | 34.14 | 34.45 | 30.14 | 0.68 | 9.37 | 1.34 |
| 600 | 2.32 | 34.77 | PPISN | 34.12 | 34.42 | 30.15 | 0.63 | 9.38 | 1.34 |